Week of 160104

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:
  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes

ALICE ATLAS CMS LHCB
  BNL FNAL  

Monday

Attendance:

  • local: Maria Alandes (chair&minutes), Maarten Litmaath (ALICE), Juan Manuel Guijarro (Grid&Batch)
  • remote: Daniele Bonacorsi (CMS), Vladimir Romanovskiy (LHCb), Michael Ernst (BNL), Lisa Giachetti (FNAL), Rolf Rumler (IN2P3), Dmytro Karpenko (NDGF), Onno Zweers (NL-T1), Kyle Gross (OSG), John Kelly (RAL), Di Qing (TRIUMF)

Experiments round table:

  • ATLAS reports (raw view) -
    • Smooth running jobs over the whole Xmas break.

  • CMS reports (raw view) -
    • Our best wishes for a great 2016 to everyone!
    • Activity was moderate/high through the break. Many thanks to all sites for their work and support.
    • Quite smooth:
      • good use of distributed resources for production campaigns
      • digi-reco in preparation for Moriond, plus data rereco, continued through the break
      • HLT usage for offline processing helped and contributed to the general efforts
    • Few issues, reported in GGUS tickets, nothing major (not a full list today)

  • ALICE -
    • Best wishes for 2016!
    • Normal to high activity levels during the break
      • Thanks to the sites for keeping things in good shape!
      • The first round of the heavy-ion reconstruction finished!
    • Dec 18: alarm ticket GGUS:118443 because the CASTOR transfer manager was stuck
      • Fixed later that afternoon, thanks!
    • Dec 31: team ticket GGUS:118554 for CASTOR because of same issue
      • OK again since Jan 1 00:00, thanks!
    • Dec 31: team ticket GGUS:118559 for EOS at CERN
      • Partly due to EOS-ALICE being ~full!
      • Some disk servers were unavailable
      • Mitigated by the admins, thanks!
    • Dec 31: tape SE at KIT working again, thanks!

  • LHCb reports (raw view) -
    • Data Processing
      • Monte Carlo running, with intention of continuing to run this during YETS (including on HLT farm); user analysis at T0/1/2D sites
      • Data prestaging for Stripping 24 Run II data is almost finished.
    • T0:
      • CREAM CEs are unavailable (GGUS:118588).
      • 50 Scheduled pilots at each CE. Some of them submitted day ago.
    • T1:
      • Staging problem at PIC,Gridka and RRCKI during holidays; Already fixed.
      • RAL: Disk server unavailable

Sites / Services round table:

  • ASGC: NA
  • BNL: NTR
  • CNAF: NA
  • FNAL: NTR
  • GridPP: NA
  • IN2P3: NTR
  • JINR: NA
  • KISTI: NA
  • KIT: NA
  • NDGF: NTR
  • NL-T1: NTR
  • NRC-KI: NA
  • OSG: NTR
  • PIC: NA
  • RAL: A disk server with ~30K files is temporarily unavailable, presumably until tomorrow afternoon.
  • TRIUMF: NTR

  • CERN batch and grid services: Juan Manuel will follow up on the CREAM CE issues reported by LHCb.
  • CERN storage services: NA
  • Databases: NA
  • GGUS: NA
  • Grid Monitoring: NA
  • MW Officer: NA

AOB:

Thursday

Attendance:

  • local: Maria Alandes (chair and minutes), Andrea Manzi (MW Officer), Xavier Espinal (Storage), Maarten Litmaath (ALICE)
  • remote: Dario Barberis (ATLAS), Daniele Bonacorsi (CMS), Zoltan (LHCb), Michel Ernst (BNL), Jeremy Coles (GridPP), Rolf Rumler (IN2P3), Christian (NDGF), Andrew Pickford (NL-T1), Rob Quick (OSG), Pepe Flix (pic), Gareth Smith (RAL), Di Qing (TRIUMF), Javier Sanchez (ATLAS-T2-ES)

Experiments round table:

  • ATLAS reports (raw view) -
    • All OK. Production continuing full steam. Reprocessing almost completed.
    • Ongoing staging errors at INFN-T1.

  • CMS reports (raw view) -
    • updated the ticket about provisioning fresh VMs for the Tier-0 processing (GGUS:118546).
      • Jan-4: all VMs clean-up up, CMS has constant job pressure, so the system just kept creating more (up to the quota limit). Jan-5: close to the quota, all successful, then some started to fail. Anyway we reached >1500 running VM instances, not having the remaining 100 or so does not change much. Jan-7: some again in error state and being cleaned-up, but overall the resource utilization is good and T0 Ops continue.
    • glexec errors at T1_DE_KIT (GGUS:118660), quick reply, seems intermittent, no obvious reasons, being kept monitored.

  • ALICE -
    • CERN: team ticket GGUS:118619 opened for CASTOR on Jan 5
      • the transfer manager was again stuck
      • debugged live by the devs
      • root cause was not found yet
      • it is now running with more debug info enabled, thanks!

  • LHCb reports (raw view) -
    • Data Processing
      • Monte Carlo and user analysis at T0/1/2D sites
      • Pre-staging the data for re-stripping is almost finished.
    • Issues
      • T1:
        • RRCKI: Staging problem
        • IN2P3: Files can not be opened, because SRM returned a wrong end point (GGUS:118655)
        • CNAF: Transfer failure (GGUS:118665)

Sites / Services round table:

  • ASGC: NA
  • BNL: NTR
  • CNAF: Salvatore Tupputi sent the report after the meeting: LHCb ticket is in progress. ATLAS ticket: the file staging has returned to normal and investigations on causes is ongoing.
  • FNAL: NA
  • GridPP: NTR
  • IN2P3: NTR
  • JINR: NA
  • KISTI: NA
  • KIT: NA
  • NDGF: NTR
  • NL-T1: NTR
  • NRC-KI: NA
  • OSG: Rob reports that Vanderbilt is getting a lot of tickets due to gLexec transient errors. Daniele explains that this could be related to the fact that they were in downtime but they declared the downtime only for part of the services and not for all of them. Daniele already notified the operators so no more tickets were opened to those services that were actually also in downtime.
  • PIC: NTR
  • RAL: A network link was down between Tuesday at 4 PM and Wednesday at 9 AM.
  • TRIUMF: NTR

  • CERN batch and grid services: NA
  • CERN storage services: Xavier reports about a couple of issues affecting EOS ALICE: on Tuesday a preventive intervention had to be made restarting the namespace. This took 40min in total. Today the instance was stuck again. It is probably related to the fact that it's full but it's still being investigated. ALICE is already doing some cleaning.
  • Databases: NA
  • GGUS: NA
  • Grid Monitoring: NA
  • MW Officer: NTR

AOB:

Edit | Attach | Watch | Print version | History: r9 < r8 < r7 < r6 < r5 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r9 - 2016-01-07 - MaartenLitmaath
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback